OpenStudy (anonymous):

What is y = x^2 -16x + 40 written in vertex form?

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

Do you know how to complete the square?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

\[y=x ^{2}-16x+40\] Do you know how to do difference of squares?

OpenStudy (anonymous):

no ma'am :/

OpenStudy (anonymous):

the equation is in standard form, therefore, it is y=x^2+bx+c divide your b value by two and square it

OpenStudy (anonymous):

oh so that means y = (x - 8)^2 + 48

OpenStudy (anonymous):

No, once you have your value, square it.

OpenStudy (anonymous):

You should get 64

OpenStudy (anonymous):

oh so ... y = (x - 16)^2 + 64 ?

OpenStudy (anonymous):

y= (x^2-16x+64)-64+40

OpenStudy (anonymous):